OpenStudy (anonymous):

How do I find the Total Curvature, of a parametrized surface with Gaussian curvature: -4/((1+u^2+v^2)^4) ?

OpenStudy (anonymous):

It maybe easier to visualise it like this: \[\int\limits_{0}^{2\pi}\int\limits_{0}^{2\pi} -\frac{ 4 }{ (1+u^2+v^2)^4 }du dv\] if anyone can help with double integrals, i think that is what the limits would be, and it's definitely the correct formula for Gaussian curvature. @agent0smith @dan815 @satellite73 @skullpatrol @thomaster I would really appreciate it if you could take a look? (I may be doing it completely wrong!) (and sorry for tagging people!)

OpenStudy (amistre64):

i just do not have enough exposure or experience with curvature, much less a gaussian curvature to be of any use with this :/

OpenStudy (anonymous):

Okay no problem :-) thanks for giving it a go anyway and taking the time to look!

OpenStudy (amistre64):

and tagging people is fine; useless at the moment, but fine. When there are no smartscores showing up the site is pretty much broken

OpenStudy (anonymous):

Oh really! I wondered what was going on with that! Okay, thanks for letting me know!

OpenStudy (dan815):

im here u still need help here?

OpenStudy (anonymous):

Yes I would love some help please!